OpenStudy (anonymous):

Algebra 1 help :D Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2x - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (darkbluechocobo):

Dinkleburg.....

OpenStudy (darkbluechocobo):

o its the same thing you were working with

OpenStudy (anonymous):

y-6=1/2(x-2) .. The rest i dont really know

OpenStudy (darkbluechocobo):

@leon88

OpenStudy (darkbluechocobo):

@chmvijay Assistance

OpenStudy (darkbluechocobo):

@iambatman Assistance Batman Assistance

OpenStudy (anonymous):

Ok so first, what's an equation of a line?

OpenStudy (anonymous):

o.o ummmm ..

OpenStudy (darkbluechocobo):

mx+b

OpenStudy (anonymous):

It's the equation y = mx + b form where m is the slope.

OpenStudy (anonymous):

the slope is going to be 2x

OpenStudy (anonymous):

lol howd you get the slope?

OpenStudy (anonymous):

Just assume because it has "x" by it

OpenStudy (anonymous):

Oh I thought you were just given the points but there is more :o

OpenStudy (anonymous):

it's the first one

OpenStudy (anonymous):

since the slope has to be the same for them to be parallel

OpenStudy (anonymous):

You can check by using the point slope form though

OpenStudy (anonymous):

Thank you :D !